FEL Upgrade Project Weekly Brief – September 2-6, 2002




The drive laser optical transport cans were rehung beginning preparation to applying drive beam to the cathode to start commissioning. They still need final alignment and beam tubes to be added.

We got shipment of a sizeable number of dipoles for the machine and established a procedure for setting accurate field levels in the major dipoles.


We hosted a visit for AES engineers for three days this week to work over design issues on the AES 750 MHz injector cavities. Substantial progress was made in moving the system engineering detailing forward toward finalization.

We hosted a visit for a group from Boeing to discuss the evolution of our FEL program. They have long term interest in being systems integrator for the Navy drawing on their experience with both FELs and other laser programs. We also identified a few areas of potential application interest.

Many in the group prepared papers and posters to be presented next week at the International FEL Conference held this year at Argonne.

We also verified that our present EA approval covers operation of the IR machine at design levels as well as the UV machine. An updated Facility Safety Assessment Document with new safety and operation envelopes for the Upgrade machine is in preparation.

WBS 4 (Injector):

All the hardware for the ball cathode has been silver-plated and about 70% of its components are ready for assembly. Some rust spots were found on the shield door mechanism during the cleaning process. It was also found that this piece became slightly magnetic. A new piece is being fabricated with 316L stainless steel. The light box has been partially cleaned and all its internal components silver-plated. The resistor drawings have been submitted to the machine shop. Polishing of the high voltage ring is in progress. The stalk tube is being smoothed out with sandpaper in preparation for vacuum.

Field emission scans were performed on two stainless steel one-inch discs polished down to one-micron finish. The results show ten emitters in the range of 80 to 140 MV/m for the bare stainless steel sample, while only two emitters were found around 120 MV/m on the ion-implanted sample. The maximum field in the cathode support tube, which has already been implanted will be around 12 MV/m.

Gun HVPS - W. Sommer has joined us for the installation and testing of this HVPS. The PSS/MPS interface to the HVPS was roughed out this morning with K. Jordan. Wiring is underway. The pressure switch, gauge, and transducer were mounted today. Documentation was received this week from Glassman for the wiring and connections for the HVPS. The final part for the mechanism combing the Running and Conditioning Resistors was signed off this week.

WBS 5 (SRF):

Construction of the cavities for the third FEL cryomodule is fully underway:

WBS 6 (RF):

Quarter Klystrons - Work continued on the unit 3 LCW manifold for the RF Circulator, RF Load, and the HOM Filter. The assembly is nearly finished, but it still has minor leaks. The R. Taylor is going to fix the leaks. Nearly all of the RF couplers were tested and verified this week.

Quarter HVPS - Unit 3 is complete and ready to power the klystron. Installation work has shifted to unit 4. Interlock tests are scheduled for early next week. The interface software to EPICS effort has started.

WBS 8 (Instrumentation):

The mechanical design for the recirculation dump (4") Beam Position Monitor (BPM) is complete and passed on to the mechanical designers. This new design will feature a ISI Seal SMA feedthrough on a mini-conflat instead of the welded feedthroughs that were prone to failure during assembly.

The design for the Personal Safety System (PSS) interface to the new gun high voltage power supply is complete. The new interface chassis is progressing well, this new design will greatly simplify the operation and trouble shooting of any high voltage problems.

The Linux frame grabber software is progressing nicely, Zafir from the operations group is also interested and may help out. Further work has been done in writing code to find the reference marks, and measuring the spot size. We have also received an email from Dean (one of our student that is back in school), and he's still interested in the project, so we'll try to keep him in the loop in case he has any suggestions or improvements.

The schematic for the Acromag fanout module has been given to the Fast Electronics (Physics - THANKS) group for layout. This will be used both for high voltage power supply control and for the drive laser control. The checkplot and schematic for the isolation module fan out board have been received from EECAD and are ready for sign off. Cables were pulled from the stepper motors chassis rack to the Optical Cavity and Lab 2. The software group has started updating the magnet controls for the 0F region using the RS-485 communications upgrade. This should be ready in about 2 weeks.

BLMs - The development of the F0151 prototype is proceeding well. All of the surface-mount components have been placed and the board is being finished in Lab 5. The F0150 board has been sent out for production. DLPC - The VME based CO305 carrier board is under final check before being sent out for production. MPS - Documentation for the MPS upgrade and preparation for the FEL conference has been the dominant priority. MISC - A discussion of the drive laser's over-all controls led to a short punch list of "to-do" items necessary to declare the drive laser "ready".

WBS 9 (Transport):


Optical Chicane Dipoles (GW)

MEI discovered the reason behind the gap height problems with the second batch of cores and is having the return legs touched up at the grinders. The coils are finished. We await the successful core check out and glue down of the Purcell gap materials and we will have a good

magnet batch.

Injector Dipoles (DU/DV)

• Success! The measurements of the bulk field on the first GV continued, finally getting repeatability after going between energies in the part in 10^5 level. Field flatness is within 2 parts in 10^4. An error in overdriving the magnet core lead us to use the trim coils to erase the magnetic memory of the magnet in one stroke rather than go through a 5 hour degaussing procedure. That led us to using a similar procedure between all magnet settings. It works beautifully.

• We are now in a position to adjust the field clamps to obtain the right field integral level and transverse field integral gradient. (Earlier measurements had proven the gradient was sufficiently linear with distance)

• Wang NMR shipped the remaining 4 GVs and 1 GU to complete the order. We now have 12 of the 22 dipoles we need for the IR machine in house.

• We are preparing the second GV for measurement as soon as the first is through with its characterization.

Arc 180 Degree Dipoles (GY)

• GY Coil Potting fixture continues its preparation for potting.

• We resolved our contractual issues with Wang NMR over winding the replacement for the first coil.

Arc Bend, Reverse Bend Dipoles (GQ, GX)

• The GX cores are in final assembly. One glitch in the cover bolt pattern became evident and is being fixed by PECo. We have had very few glitches of this sort ? a tribute to the quality of the drafting and the use of relational 3 dimensional drafting models.

• At Wang NMR, potting fixtures are in final preparation for the GQ and GX coils.


Trim Quad (QT)

• We started formal measurement of four trim quads on the test station after resolving the problem with the test stand. The temperature problem with the side coils has been fixed with the new cooling plates.

Sextupole (SF)

• DULY Research should finish all drawings today, making them ready for signature.

• We have decided to fore-go immediate fabrication of the SF sextupoles for eight locations in favor of using modified and doubled-up IR Demo Sextupoles in two locations in the 2nd Arc. David Douglas believes this configuration will still allow lasing, but at a small reduced intensity during a commissioning period. This grace period will allow time for fabrication of the SF sextupoles. They can then be installed later to gain full intensity.

Octupole (OT)

• The Octupole design is about half complete at DULY Research while they finish the Sextupole.

Beam Line and Vacuum

• X-Chamber - The third transition (with small machining gouge) was approved for use in its chamber. Final rework and machining on the x-chamber and transitions (for weld preparations) is in progress. Welding is anticipated to begin next week.

• Chicane Chamber - Wire EDM work for the transition piece has been quoted. Chamber plates, tubes, and flanges are on hold to allow higher priority parts to be fabricated.

• Y-Chamber - All material for the Y chamber has been received. Plate machining will begin next week.

• ARC Chambers - Master Machine is f putting together and welding the first chamber. All the flanges have been machined. The last transition is on the wire EDM machine. The windows for the 180° chambers are 75% complete. All water jet cutting is complete and machining of the bevels is expected to finish by Sept. 13. They will know more after the completion of the first chamber and feel that completion by the end of September is still obtainable.

• An alternate configuration for the paddle viewer ports was approved and will be implemented into chambers using redline updates of drawing dimensions.

• Design of girders for the regions around the wiggler was suspended for higher priority work.

• The stands for sextupoles and trim quads were received.

• Girder and vacuum assembly continued

• Low Conductivity Water piping installation continued in the FEL enclosure

• The alignment crew continued stand placement.

WBS 11 (Optics):

This week's mechanical work on the optical cavity was spent primarily measuring the mechanical resonances of the 3" deformable mirror assembly. This information is being used to better model the feedback circuit required for the active mirror steering. We found a commercial coating vendor to sputter tungsten disulfide onto our leadscrews. Two leadscrews are being stripped of their silver coating and will be sent to the vendor for coating. The 2" outcoupler mirror mounts (2") arrived over two weeks early and are being prepared for the subsequent brazing and plating steps.

Detailing continues at a rapid rate on the optical transport turning cassette, the collimator and insertable mirror designs. As drawings for subassemblies are complete they are checked and sent out for procurement. We are putting together orders for other parts that will be used for these assemblies. Optical modeling shows that the additional distance between the outcoupler and the collimator does not necessitate a complete redesign of the outcouplers.

We assisted in reinstallation of optical components for the drive laser optical transport. We received spare mirrors for the transport, and they will be QA'ed.

Other activities:

We prepared presentations to be given at the FEL conference and continued work on the presentation at the Boulder Damage Conference. We did some further work (fabrication of hardware) on the JTO-funded cryomirror project.